OpenStudy (anonymous):

can someone please show all the steps to find the difference: I am totally lost (-5x^2 + 2x – 5) – (7x^2 – 2x – 3)

OpenStudy (anonymous):

Distribute the - sign to all terms on the right parenthesis and then add like terms

OpenStudy (anonymous):

first rewrite without parentheses. don't touch anything in the first parentheses, just rewrite the terms \[-5x^2+2x-5\]

OpenStudy (anonymous):

for the second one, distribute the minus sign, i.e. remover the parentheses and change all the signs \[– (7x^2 – 2x – 3) =-7x^2+2x+3\]

OpenStudy (anonymous):

\[-5x^2-7x^2 +2x+ 2x -5 +3\] Quite sure you can go on from there. I just put the like terms together for you

OpenStudy (anonymous):

this give \[(-5x^2 + 2x – 5) – (7x^2 – 2x – 3) \] \[=-5x^2 + 2x – 5– 7x^2 +2x + 3 \]

OpenStudy (anonymous):

don't skip steps or try to do this in your head. write it out on paper now combine like terms

OpenStudy (anonymous):

^^ and just practise, practice makes perfect.

OpenStudy (anonymous):

so the final answer would be 5x^2 + 2x - 5- 7x^2 + 2x + 3

OpenStudy (anonymous):

frustration, frustration frustration

OpenStudy (anonymous):

You would have to add: \[-5x^2-7x^2\] \[2x+2x\] \[-5+3\] Hint: \[2x+2x = 4x\]

OpenStudy (anonymous):

would that not be 4x^2 ? ?

OpenStudy (anonymous):

\[2x + 2x \] ?

OpenStudy (anonymous):

ok i c just the 4x

OpenStudy (anonymous):

\[2x+2x=4x \]\[2x \times 2x = 4x^2\]

OpenStudy (anonymous):

ok you are right thanks

OpenStudy (anonymous):

So now just add the other terms together and let me know what you get...That should be your answer.

OpenStudy (anonymous):

ok